# Quantifying the noise sensitivity of the Wasserstein metric for images

### v1

@@ A.2 Proof of Theorems
$(W_{p}^{\pm})^{p}(\mu,\mu+\epsilon)=W_{p}^{p}\bigg(\mu+(\mu+\epsilon)_{-},(\mu+\epsilon)_{+}\bigg).$ (35)
Thus, using Proposition 3 on $\mu+(\mu+\epsilon)_{-}$ and $(\mu+\epsilon)_{+}$, we get
$W_{p}^{p}\bigg(\mu+(\mu+\epsilon)_{-},(\mu+\epsilon)_{+}\bigg)$ (36)
$\leq W_{p}^{p}\bigg(\big(\mu+(\mu+\epsilon)_{-}-(\mu+\epsilon)_{+}\big)_{+},\big((\mu+\epsilon)_{+}-(\mu+(\mu+\epsilon)_{-})\big)_{+}\bigg)$ (37)
$=W_{p}^{p}\bigg(\big(\mu-((\mu+\epsilon)_{+}-(\mu+\epsilon)_{-})\big)_{+},\big((\mu+\epsilon)_{+}-(\mu+\epsilon)_{-}-\mu)\big)_{+}\bigg)$ (38)
$=W_{p}^{p}\bigg(\big(\mu-(\mu+\epsilon)\big)_{+},\big(\mu+\epsilon-\mu\big)_{+}\bigg)$ (39)
$=W_{p}^{p}\bigg((-\epsilon)_{+},\epsilon_{+}\bigg)=W_{p}^{p}(\epsilon_{-},\epsilon_{+}).\qed$ (40)
Proof of Theorem 2.
Using the Kantorovich-Rubinsten duality,
$W^{\pm}_{1}(\mu,\mu+\varepsilon)=\sup_{f\in\mathrm{Lip}_{1}}\langle f,\varepsilon\rangle=W_{1}(\varepsilon^{+},\varepsilon^{-}).$ (41)
The first equality is the signed dual form with $\mu-(\mu+\varepsilon)=-\varepsilon$. For the second, $\int\varepsilon=0$ implies $\varepsilon=\varepsilon^{+}-\varepsilon^{-}$ with equal masses, so the balanced duality gives $W_{1}(\varepsilon^{+},\varepsilon^{-})=\sup_{f\in Lip_{1}}\langle f,\varepsilon\rangle$
Let $m=\varepsilon^{+}(G_{n})=\varepsilon^{-}(G_{n})$. By homogeneity of $W_{1}$,
$W_{1}(\varepsilon^{+},\varepsilon^{-})=m\,W_{1}\Big(\tfrac{\varepsilon^{+}}{m},\tfrac{\varepsilon^{-}}{m}\Big).$ (42)
Apply Proposition 1 to the probability measures $\varepsilon^{+}/m$ and $\varepsilon^{-}/m$. There exists an integer $k^{*}$ with $k^{*}\asymp\log_{2}n$ such that
$W_{1}(\frac{\varepsilon^{+}}{m},\frac{\varepsilon^{-}}{m})\leq\frac{\sqrt{2}}{2}\,2^{-k^{*}}+\frac{\sqrt{2}}{2}\sum_{k=0}^{k^{*}}2^{-k}\sum_{Q\in\mathcal{D}_{k}}|(\frac{\varepsilon^{+}}{m}-\frac{\varepsilon^{-}}{m})(Q)|.$ (43)
Multiplying by $m$ gives
$W_{1}(\varepsilon^{+},\varepsilon^{-})\leq\frac{\sqrt{2}}{2}\,m\,2^{-k^{*}}+\frac{\sqrt{2}}{2}\sum_{k=0}^{k^{*}}2^{-k}\sum_{Q\in\mathcal{D}_{k}}\Big|\sum_{x\in Q}\varepsilon(x)\Big|.$ (44)
$\mathbb{E}W_{1}(\varepsilon^{+},\varepsilon^{-})\leq\frac{\sqrt{2}}{2}\,2^{-k^{*}}\,\mathbb{E}m+\frac{\sqrt{2}}{2}\sum_{k=0}^{k^{*}}2^{-k}\sum_{Q\in\mathcal{D}_{k}}\mathbb{E}\Big|\sum_{x\in Q}\varepsilon(x)\Big|.$ (45)
Since each $\varepsilon(x)$ is Gaussian with variance $\sigma^{2}$, one has $\mathbb{E}|\sum_{x\in Q}\varepsilon(x)|\leq\sigma\sqrt{|Q|}\sqrt{2/\pi}$ as well as $\mathbb{E}m=\sum_{x\in G_{n}}\mathbb{E}(\varepsilon(x))_{+}=n^{2}\sigma/\sqrt{2\pi}$. Furthermore, the dyadic family $\mathcal{D}_{k}$ has $|\mathcal{D}_{k}|=2^{2k}$ cubes of cardinality $|Q|=n^{2}/2^{2k}$. Therefore
Plugging this into the multiscale sum yields
$\frac{\sqrt{2}}{2}\sum_{k=0}^{k^{*}}2^{-k}\sum_{Q\in\mathcal{D}_{k}}\mathbb{E}\Big|\sum_{x\in Q}\varepsilon(x)\Big|\leq\frac{\sqrt{2}}{2}\sigma\sqrt{\tfrac{2}{\pi}}\,n\sum_{k=0}^{k^{*}}1\leq\frac{\sqrt{2}}{2}\sigma\sqrt{\tfrac{2}{\pi}}\,n\big(k^{*}+1\big).$ (47)
With $k^{*}\asymp\log_{2}n$ this gives the $\sigma n\log_{2}n$ contribution.
For the coarse term choose $k^{*}$ so that $2^{-k^{*}}\asymp 1/n$. Then
$\frac{\sqrt{2}}{2}\,2^{-k^{*}}\,\mathbb{E}m\asymp\frac{\sqrt{2}}{2}\frac{1}{n}\cdot\frac{n^{2}\sigma}{\sqrt{2\pi}}=\frac{\sqrt{2}}{2}\frac{\sigma n}{\sqrt{2\pi}},$ (48)
which is the $\sigma n$ contribution.
$\mathbb{E}W_{1}(\varepsilon^{+},\varepsilon^{-})\leq\frac{2}{\sqrt{\pi}}\sigma n\log_{2}n+\frac{1}{2\sqrt{\pi}}\sigma n.$ (49)
In this derivation the factor $m$ appears only in the coarse term and contributes to the $\sigma n$ piece after expectation. In the oscillation terms it cancels with the normalization, so no additional dependence on $m$ remains. There is no additive grid term independent of $\sigma$, hence no $1/(\sqrt{2}n)$ tail. ∎
Proof of Theorem 3.
By Proposition 5, we only need to upper bound $W_{p}({\epsilon}_{+},{\epsilon}_{-})$.
As, by assumption the noise has total zero mass, this quantity is well defined.
Then, by the multiscale bound of Proposition 1
$W_{p}^{p}({\epsilon}_{+},{\epsilon}_{-})$ $=2^{-p/2}{\epsilon}_{+}(G_{n})W_{p}^{p}\left(\frac{{\epsilon}_{+}}{{\epsilon}_{+}(G_{n})},\frac{{\epsilon}_{-}}{{\epsilon}_{+}(G_{n})}\right)$ (50)
$\leq 2^{-pk^{*}-p/2}{\epsilon}_{+}(G_{n})+2^{-p/2}\sum_{k=1}^{k^{*}}2^{-p(k-1)}\sum_{Q_{i}^{k}\in\mathcal{Q}^{k}}\lvert{\epsilon}_{+}(Q_{i}^{k})-{\epsilon}_{-}(Q_{i}^{k})\rvert$ (51)
$\leq 2^{-pk^{*}-1/2}{\epsilon}_{+}(G_{n})+2^{-p/2}\sum_{k=1}^{k^{*}}2^{-p(k-1)}\sum_{Q_{i}^{k}\in\mathcal{Q}^{k}}\lvert{\epsilon}(Q_{i}^{k})\rvert$. (52)
$\mathbb{E}\sum_{Q\in\mathcal{Q}_{k}}|\epsilon^{*}(Q)|\leq 4^{k}\sqrt{\frac{2}{\pi}}\sigma 2^{\eta-k}.$ (53)
As in the previous proof,
$\mathbb{E}{\epsilon}_{+}(G_{n})=\frac{n^{2}}{2}\sqrt{\frac{2}{\pi}}\sigma\sqrt{1-\frac{1}{n^{2}}}.$ (54)
Altogether,
$\mathbb{E}W_{p}^{p}({\epsilon}_{+},{\epsilon}_{-})$ $\leq 2^{-pk^{*}-1/2}\frac{4^{\eta}}{2}\sqrt{\frac{2}{\pi}}\sigma+2^{\eta}2^{p/2}\sum_{k=1}^{k^{*}}2^{-(p-1)k}\sqrt{\frac{2}{\pi}}\sigma$ (55)
$\leq 2^{-pk^{*}-1/2}\frac{4^{\eta}}{2}\sqrt{\frac{2}{\pi}}\sigma+2^{\eta}2^{p/2}\sqrt{\frac{2}{\pi}}\sigma\ \frac{1-2^{-(p-1)k^{*}}}{2^{p-1}-1}.$ (56)
We take $k^{*}=\eta$ again to get
$\mathbb{E}W_{p}^{p}({\epsilon}_{+},{\epsilon}_{-})$ $\leq 2^{-(p-1)\eta}2^{-1/2}\frac{2^{\eta}}{2}\frac{1}{\sqrt{\pi}}\sigma+2^{\eta}2^{p/2}\sqrt{\frac{2}{\pi}}\sigma\ \frac{1}{2^{p-1}-3/2}$ (58)
$\leq\frac{2^{\eta}}{\sqrt{\pi}}\sigma\left(2^{-(p-1)\eta-3/2}+\frac{2^{(p+1)/2}}{2^{p-1}-1}\right).$ (59)
Remark that $2^{-(p-1)\eta-3/2}\leq 1/2$ and that $\frac{2^{(p+1)/2}}{2^{p-1}-1}$ is decreasing with value $2\sqrt{2}$ at 2. The claim follows. ∎
@@
For any two images $\mu,\nu:G_{n}\to[0,\infty)$ and independent noises $\epsilon_{\mu},\epsilon_{\nu}$ as in Assumption 1,
$W_{1}([\mu+\epsilon_{\mu}-\nu-\epsilon_{\nu}]_{+},[\nu+\epsilon_{\nu}-\mu-\epsilon_{\mu}]_{+})$
$\qquad\leq W_{1}(\mu,\nu)+W_{1}\big((\epsilon_{\mu}-\epsilon_{\nu})^{+},(\epsilon_{\mu}-\epsilon_{\nu})^{-}\big).$
@@
$W_{1}\left([\mu+\epsilon_{\mu}-\nu-\epsilon_{\nu}]_{+},[\nu+\epsilon_{\nu}-\mu-\epsilon_{\mu}]_{+}\right)=\sup_{\|f\|_{\mathrm{Lip}}\leq 1}\int f(\mu-\nu)+\int f(\epsilon_{\mu}-\epsilon_{\nu}).$ (60)
@@
$\sup_{\|f\|_{\mathrm{Lip}}\leq 1}\int f(\epsilon_{\mu}-\epsilon_{\nu})\leq W_{1}((\epsilon_{\mu}-\epsilon_{\nu})^{+},(\epsilon_{\mu}-\epsilon_{\nu})^{-})$ (62)
@@
Proof of Theorem 4.
Recall that $W_{1}^{\pm}$ satisfies the triangle inequality, so
$W_{1}^{\pm}(\mu+{\epsilon}_{\mu},\nu+{\epsilon}_{\nu})\leq W_{1}^{\pm}(\mu+{\epsilon}_{\mu},\mu)+W_{1}^{\pm}(\mu,\nu)+W_{1}^{\pm}(\nu,\nu+{\epsilon}_{\nu}).$ (63)
By symmetry
$\mathbb{E}W_{1}^{\pm}(\mu+{\epsilon}_{\mu},\mu)=\mathbb{E}W_{1}^{\pm}(\nu,\nu+{\epsilon}_{\nu})$ (64)
Therefore,
$\mathbb{E}[W_{1}^{\pm}(\mu+{\epsilon}_{\mu},\nu+{\epsilon}_{\nu})-W_{1}^{\pm}(\mu,\nu)]\leq 2\mathbb{E}W_{1}^{\pm}(\mu,\mu+{\epsilon}_{\mu}).$ (65)
$W_{1}^{\pm}(\mu,\mu+{\epsilon})$ $=W_{1}(\mu_{+}+(\mu+{\epsilon})_{-},(\mu+{\epsilon})_{+}+\mu_{-})$ (66)
$=W_{1}(\mu+(\mu+{\epsilon})_{-},(\mu+{\epsilon})_{+})$ $\text{(since $\mu_{+}=\mu$ and $\mu_{-}=0$)}.$ (67)
$W_{1}^{\pm}$ $\left(\mu,\mu_{\epsilon}\right)=(\mu+{\epsilon})_{+}(G_{n})W_{1}^{\pm}\left(\frac{\mu+(\mu+{\epsilon})_{-}}{(\mu+{\epsilon})_{+}(G_{n})},\frac{(\mu+{\epsilon})_{+}}{(\mu+{\epsilon})_{+}(G_{n})}\right)$ (68)
$\leq\tfrac{\sqrt{2}}{2}\cdot 2^{-k^{*}}(\mu+{\epsilon})_{+}(G_{n})+\tfrac{\sqrt{2}}{2}\sum_{k=1}^{k^{*}}2^{-(k-1)}\sum_{Q_{i}^{k}\in\mathcal{Q}^{k}}\big\lvert(\mu+(\mu+{\epsilon})_{-})(Q_{i}^{k})-(\mu+{\epsilon})_{+}(Q_{i}^{k})\big\rvert$.
By considering the two cases $(\mu+{\epsilon})(Q_{i}^{k})\geq 0$ and $(\mu+{\epsilon})(Q_{i}^{k})<0$ it is easy to see that the term $(\mu+(\mu+{\epsilon})_{-})(Q_{i}^{k})-(\mu+{\epsilon})_{+}(Q_{i}^{k})$ is equal to ${\epsilon}(Q_{i}^{k})$, so the bound above simplifies to
$W_{1}^{\pm}\left(\mu,\mu_{\epsilon}\right)$ $\leq 2^{-k^{*}-\tfrac{1}{2}}(\mu+{\epsilon})_{+}(G_{n})+\frac{\sqrt{2}}{2}\sum_{k=1}^{k^{*}}2^{-(k-1)}\sum_{Q_{i}^{k}\in\mathcal{Q}^{k}}\lvert{\epsilon}(Q_{i}^{k})\rvert$. (69)
Rewrite the noise as ${\epsilon}={\epsilon}^{\prime}-\bar{\epsilon}$ where $\epsilon$$^{\prime}$ is i.i.d. $\mathcal{N}(0,\sigma^{2})$ at each pixel and $\bar{\epsilon}\in\mathbb{R}$ is the mean of all $\epsilon^{\prime}$ terms across the entire image. Since $Q_{i}^{k}$ is a square region of size $2^{\eta-k}\times 2^{\eta-k}$ and $\bar{{\epsilon}}$ is the mean of $4^{\eta}$ i.i.d. Gaussian noise terms, it follows that ${\epsilon}^{\prime}(Q_{i}^{k})\sim\mathcal{N}(0,4^{\eta-k}\sigma^{2})$ and also that $\bar{{\epsilon}}\sim\mathcal{N}(0,\sigma^{2}/4^{\eta})=\mathcal{N}(0,\sigma^{2}/n^{2})$. Recall that $\mathbb{E}\lvert X\rvert=\sigma\sqrt{2/\pi}$ when $X\sim\mathcal{N}(0,\sigma^{2})$.
Since $\epsilon^{*}(Q_{i}^{k})=\sum_{x\in Q_{i}^{k}}\epsilon^{\prime}(x)-4^{\eta-k}\bar{\epsilon}$,
$\operatorname{Var}\big(\epsilon^{*}(Q_{i}^{k})\big)=\sigma^{2}\Big(4^{\eta-k}+\frac{4^{2(\eta-k)}}{n^{2}}-2\frac{4^{2(\eta-k)}}{n^{2}}\Big).$ (70)
Thus,
$\mathbb{E}|\epsilon^{*}(Q_{i}^{k})|=\sqrt{\frac{2}{\pi}}\sigma 2^{\eta-k}\Big(1-n^{2}4^{-k}\Big)^{1/2}.$ (71)
Summing over the $4^{k}$ cells at level $k$,
$\mathbb{E}\sum_{Q\in\mathcal{Q}_{k}}|\epsilon^{*}(Q)|=4^{k}\sqrt{\frac{2}{\pi}}\sigma 2^{\eta-k}\Big(1-n^{2}4^{-k}\Big)^{1/2}.$ (72)
$\mathbb{E}\left[\frac{\sqrt{2}}{2}\sum_{k=1}^{k^{*}}2^{-(k-1)}\sum_{Q_{i}^{k}\in\mathcal{Q}^{k}}\lvert{\epsilon}(Q_{i}^{k})\rvert\right$] $\leq\frac{\sqrt{2}}{2}\sum_{k=1}^{k^{*}}2^{-(k-1)}4^{k}\sqrt{\frac{2}{\pi}}\sigma 2^{\eta-k}$ (73)
$=\frac{2^{\eta+1}\sigma}{\sqrt{\pi}}k^{*}.$ (74)
We take $k^{*}=\eta=\log_{2}n$ to obtain the bound
$\mathbb{E}W_{1}^{\pm}\left(\mu,\mu_{\epsilon}\right)\leq\frac{1}{\sqrt{2}n}\mathbb{E}\left[(\mu+{\epsilon})_{+}(G_{n})\right]+\frac{2n\log_{2}n}{\sqrt{\pi}}\sigma.$ (75)
We now bound the first term in the RHS.
$\mathbb{E}\left[(\mu+{\epsilon})_{+}(G_{n})\right]\leq\mathbb{E}[\mu_{+}(G_{n})]$ $+\mathbb{E}[{\epsilon}_{+}(G_{n})]$ (76)
$=1$ $+\mathbb{E}[{\epsilon}_{+}(G_{n})]$ (77)
$\mathbb{E}{\epsilon}_{+}(G_{n})=\tfrac{1}{2}\mathbb{E}|{\epsilon}|(G_{n}).$ (78)
Further set $m=\frac{1}{2}\sum_{x\in G_{n}}|\epsilon(x)|$ and recall that $\epsilon(x)\sim\mathcal{N}\big(0,\sigma^{2}(1-1/n^{2})\big)$ to derive
$\mathbb{E}m=\frac{n^{2}}{2}\sqrt{\frac{2}{\pi}}\sigma\sqrt{1-\frac{1}{n^{2}}}.$ (79)
Thus,
$\frac{1}{\sqrt{2}n}\mathbb{E}\left[(\mu+{\epsilon})_{+}(G_{n})\right]\leq\frac{1}{\sqrt{2}n}+\frac{\sigma}{2\sqrt{\pi}}n.$ (80)
Plugging this back into equation 75 gives
$\mathbb{E}W_{1}^{\pm}\left(\mu,\mu_{\epsilon}\right)\leq\frac{2n\log_{2}n+n/2}{\sqrt{\pi}}\sigma+\frac{1}{\sqrt{2}n}.$ (81)
Note that the same bound applies to $\mathbb{E}W_{1}^{\pm}\left(\nu,\nu+{\epsilon}_{\nu}\right)$. By subtracting $W_{1}^{\pm}(\mu,\nu)$ from both sides of equation 63 and taking expectations, we have
$\mathbb{E}\left[W_{1}^{\pm}(\mu_{\epsilon},\nu_{\epsilon})-W_{1}^{\pm}(\mu,\nu)\right]$ $\leq\mathbb{E}W_{1}^{\pm}(\mu_{\epsilon},\mu)+\mathbb{E}W_{1}^{\pm}(\nu,\nu_{\epsilon})$
$\leq\frac{4n\log_{2}n+n}{\sqrt{\pi}}\sigma+\frac{\sqrt{2}}{n}.\qed$ (82)
Proof of Theorem 5.
Using Proposition  4
$\mathbb{E}\big[W_{p}^{\pm}(\mu+\epsilon_{\mu},\nu+\epsilon_{\nu})\big]\leq\mathbb{E}\bigg[D^{1-\frac{1}{p}}\big(W_{1}(\mu,\nu)+W_{1}(\epsilon^{*}_{+},\epsilon^{*}_{-})\big)^{\frac{1}{p}}\bigg]$ (83)
The function $t\mapsto t^{1/p}$ is concave on $0,\infty)$, hence by Jensen:
$\mathbb{E}\bigg[D^{1-\frac{1}{p}}\big(W_{1}(\mu,\nu)+W_{1}(\epsilon^{*}_{+},\epsilon^{*}_{-})\big)^{\frac{1}{p}}\bigg]\leq D^{1-\frac{1}{p}}\bigg(\mathbb{E}\big[W_{1}(\mu,\nu)\ +W_{1}(\epsilon^{*}_{+},\epsilon^{*}_{-})\big]\bigg)^{\frac{1}{p}}$ (84)
By the linearity of expectation,
$D^{1-\frac{1}{p}}\bigg(\mathbb{E}\big[W_{1}(\mu,\nu)\ +W_{1}(\epsilon^{*}_{+},\epsilon^{*}_{-})\big]\bigg)^{\frac{1}{p}}=D^{1-\frac{1}{p}}\bigg(W_{1}(\mu,\nu)\ +\mathbb{E}\big[W_{1}(\epsilon^{*}_{+},\epsilon^{*}_{-})\big]\bigg)^{\frac{1}{p}}$ (85)
Finally, using Theorem  2 we get that
$D^{1-\frac{1}{p}}\bigg(W_{1}(\mu,\nu)\ +\mathbb{E}\big[W_{1}(\epsilon^{*}_{+},\epsilon^{*}_{-})\big]\bigg)^{\frac{1}{p}}\leq D^{1-\frac{1}{p}}\bigg(W_{1}(\mu,\nu)\ +\frac{2\sqrt{2}}{\sqrt{\pi}}\sigma n\log_{2}n+\sqrt{\frac{2}{\pi}}\sigma n.\bigg)^{\frac{1}{p}}$ (86)
Using Jensen,
$\mathbb{E}\big[W_{p}^{\pm}(\mu+\epsilon_{\mu},\nu+\epsilon_{\nu})\big]\leq(\frac{\sqrt{2}}{2})^{1-\frac{1}{p}}W_{1}(\mu,\nu)^{\frac{1}{p}}+\frac{\sqrt{2}}{2}\bigg(\frac{4}{\sqrt{\pi}}n\log_{2}n+\frac{2}{\sqrt{\pi}}n\bigg)^{\frac{1}{p}}\sigma^{\frac{1}{p}}.$ (87)

### v2

@@ A.1 Proof of Propositions
$(W_{p}^{\pm})^{p}(\mu,\mu+\varepsilon)=W_{p}^{p}\bigg(\mu+(\mu+\varepsilon)_{-},(\mu+\varepsilon)_{+}\bigg).$ (29)
Thus, using Proposition 3 on $\mu+(\mu+\varepsilon)_{-}$ and $(\mu+\varepsilon)_{+}$, we get
$W_{p}^{p}\bigg(\mu+(\mu+\varepsilon)_{-},(\mu+\varepsilon)_{+}\bigg)$ (30)
$\leq W_{p}^{p}\bigg(\big(\mu+(\mu+\varepsilon)_{-}-(\mu+\varepsilon)_{+}\big)_{+},\big((\mu+\varepsilon)_{+}-(\mu+(\mu+\varepsilon)_{-})\big)_{+}\bigg)$ (31)
$=W_{p}^{p}\bigg(\big(\mu-((\mu+\varepsilon)_{+}-(\mu+\varepsilon)_{-})\big)_{+},\big((\mu+\varepsilon)_{+}-(\mu+\varepsilon)_{-}-\mu)\big)_{+}\bigg)$ (32)
$=W_{p}^{p}\bigg(\big(\mu-(\mu+\varepsilon)\big)_{+},\big(\mu+\varepsilon-\mu\big)_{+}\bigg)$ (33)
$=W_{p}^{p}\bigg((-\varepsilon)_{+},\varepsilon_{+}\bigg)=W_{p}^{p}(\varepsilon_{-},\varepsilon_{+}).\qed$ (34)
@@
For any two images $\mu,\nu:G_{n}\to[0,\infty)$ and independent noises $\varepsilon_{\mu},\varepsilon_{\nu}$ as in Assumption 1,
$W_{1}([\mu+\varepsilon_{\mu}-\nu-\varepsilon_{\nu}]_{+},[\nu+\varepsilon_{\nu}-\mu-\varepsilon_{\mu}]_{+})$
$\qquad\leq W_{1}(\mu,\nu)+W_{1}\big((\varepsilon_{\mu}-\varepsilon_{\nu})_{+},(\varepsilon_{\mu}-\varepsilon_{\nu})_{-}\big).$
@@
$W_{1}\left([\mu+\varepsilon_{\mu}-\nu-\varepsilon_{\nu}]_{+},[\nu+\varepsilon_{\nu}-\mu-\varepsilon_{\mu}]_{+}\right)=\sup_{\|f\|_{\mathrm{Lip}}\leq 1}\int f(\mu-\nu)+\int f(\varepsilon_{\mu}-\varepsilon_{\nu}).$ (35)
@@
$\sup_{\|f\|_{\mathrm{Lip}}\leq 1}\int f(\varepsilon_{\mu}-\varepsilon_{\nu})\leq W_{1}((\varepsilon_{\mu}-\varepsilon_{\nu})_{+},(\varepsilon_{\mu}-\varepsilon_{\nu})_{-})$ (37)
@@
Proposition 6.
Let $\mu,\nu:G_{n}\to[0,\infty)$ be images on the square grid $G_{n}$ with spacing $h=1/n$, and let $\varepsilon_{\mu},\varepsilon_{\nu}$ satisfy Assumption 1. Identifying $G_{n}$ with the $2$-torus, let $D:=\mathrm{diam}(G_{n})=\sqrt{2}/2$. Then, for any $p\geq 1$,
$W_{p}^{\pm}(\mu+\varepsilon_{\mu},\nu+\varepsilon_{\nu})\leq D^{1-\frac{1}{p}}\big(W_{1}(\mu,\nu)+W_{1}(\varepsilon^{*}_{+},\varepsilon^{*}_{-})\big)^{\frac{1}{p}},\qquad\varepsilon^{*}:=\varepsilon_{\mu}-\varepsilon_{\nu}$ (38)
Proof.
By definition of the signed distance,
$W_{p}^{\pm}(\mu+\varepsilon_{\mu},\nu+\varepsilon_{\nu})=W_{p}\big((\mu{+}\varepsilon_{\mu})_{+}+(\nu{+}\varepsilon_{\nu})_{-},\ (\nu{+}\varepsilon_{\nu})_{+}+(\mu{+}\varepsilon_{\mu})_{-}\big).$ (39)
$W_{p}^{\pm}(\mu+\varepsilon_{\mu},\nu+\varepsilon_{\nu})\leq W_{p}\big([\mu{+}\varepsilon_{\mu}{-}\nu{-}\varepsilon_{\nu}]_{+},\ [\nu{+}\varepsilon_{\nu}{-}\mu{-}\varepsilon_{\mu}]_{+}\big).$ (40)
Applying this to $(\alpha,\beta)=([\mu{+}\varepsilon_{\mu}{-}\nu{-}\varepsilon_{\nu}]_{+},[\mu{+}\varepsilon_{\mu}{-}\nu{-}\varepsilon_{\nu}]_{-})$ yields
$W_{p}([\mu{+}\varepsilon_{\mu}{-}\nu{-}\varepsilon_{\nu}]_{+},[\mu{+}\varepsilon_{\mu}{-}\nu{-}\varepsilon_{\nu}]_{-})\leq D^{1-\frac{1}{p}}\bigg(W_{1}([\mu{+}\varepsilon_{\mu}{-}\nu{-}\varepsilon_{\nu}]_{+},[\mu{+}\varepsilon_{\mu}{-}\nu{-}\varepsilon_{\nu}]_{-})\bigg)^{\frac{1}{p}}.$ (42)
Using Proposition 5, we conclude
$D^{1-\frac{1}{p}}\bigg(W_{1}([\mu{+}\varepsilon_{\mu}{-}\nu{-}\varepsilon_{\nu}]_{+},[\mu{+}\varepsilon_{\mu}{-}\nu{-}\varepsilon_{\nu}]_{-})\bigg)^{\frac{1}{p}}$
$\qquad\leq D^{1-\frac{1}{p}}\bigg(W_{1}(\mu,\nu)+W_{1}\big((\varepsilon_{\mu}-\varepsilon_{\nu})_{+},(\varepsilon_{\mu}-\varepsilon_{\nu})_{-}\big)\bigg)^{\frac{1}{p}}.$
Since both $\varepsilon_{\mu}$ and $\varepsilon_{\nu}$ are normally distributed, we can say that $\varepsilon^{*}:=\varepsilon_{\mu}-\varepsilon_{\nu}$ is also normally distributed, with $\operatorname{cov}(\varepsilon^{*})=2\operatorname{cov}(\varepsilon_{\mu})$. Thus,
$D^{1-\frac{1}{p}}\big(W_{1}(\mu,\nu)+W_{1}((\varepsilon_{\mu}-\varepsilon_{\nu})_{+},(\varepsilon_{\mu}-\varepsilon_{\nu})_{-})\big)^{\frac{1}{p}}\leq D^{1-\frac{1}{p}}\big(W_{1}(\mu,\nu)+W_{1}(\varepsilon^{*}_{+},\varepsilon^{*}_{-})\big)^{\frac{1}{p}}.$ (43)
A.2 Proof of Theorems
Theorem  1
Consider two $n\times n$ images $\mu$ and $\nu$ having at least $\lambda n^{2}$, $\lambda\in(0,1]$ nonzero pixels. Assume that $\varepsilon_{\mu},\varepsilon_{\nu}$ are $\mathcal{N}(0_{n^{2}},\sigma^{2}I_{n^{2}})$. Recall the definition of $\bar{S}_{\mu_{\varepsilon},\nu_{\varepsilon}},\bar{T}_{\mu_{\varepsilon},\nu_{\varepsilon}}$ in equation 8. Then,
$W_{1}^{\pm}(\bar{S}_{\mu_{\varepsilon},\nu_{\varepsilon}},\bar{T}_{\mu_{\varepsilon},\nu_{\varepsilon}})=\frac{1}{\sum_{x\in G_{n}}S_{\mu_{\varepsilon},\nu_{\varepsilon}}(x)}\sup_{f\in\operatorname{Lip}_{1}}\left\langle f,S_{\mu_{\varepsilon},\nu_{\varepsilon}}-T_{\mu_{\varepsilon},\nu_{\varepsilon}}\left(1+\mathrm{O}_{p}\left(\frac{\sigma}{n}\right)\right)\right\rangle.$ (44)